# Shortest Paths in a Weighted Simplicial Complex

### v1

@@ Abstract
Simplicial complexes are extensively studied in the field of algebraic topology. They have gained attention in recent time due to their applications in fields like theoretical distributed computing and simplicial neural networks. Graphs are mono-dimensional simplicial complex. Graph theory has application in topics like theoretical computer science, operations research, bioinformatics and social sciences. This makes it natural to try to adapt graph-theoretic results for simplicial complexes, which can model more intricate and detailed structures appearing in real-world systems. In this article, we define the concept of weighted simplicial complex and $d$-path in a simplicial complex. Both these concepts have the potential to have numerous real-life applications. Next, we provide a novel algorithm to find the shortest paths in a weighted simplicial complex. The core principles of our algorithm align with those of Dijkstra’s algorithm for graphs. Hence, this work lays another brick for the sake of integrating graph-theoretic concepts with abstract simplicial complexes.
@@
Weighted Simplicial Complex, Shortest Path, Greedy Algorithm, Dijkstra’s Algorithm
@@ 1 Introduction
Now a graph can be seen as a boiled down version of a simplicial complex. While a graph is typically a 1-dimensional object, an abstract simplicial complex can exist in any finite dimension. One of the main differences between a graph and a simplicial complex is that in a simplicial complex with $d\geq 2$, more than two $((d-1))$-dimensional simplices can be connected by a single $d$-dimensional simplex, whereas in a graph, an edge can connect at most two vertices.
@@
On the other hand, Simplicial complexes have also been used in various fields as detailed in [6, 5, 7]. For example, in [6] and [5], the results on simplicial complexes were used to prove new results in theoretical distributed computing. Distributed computing is a field of computer science that studies distributed systems. In [7], Kozlov studied the chromatic subdivision of a simplicial complex. Chromatic subdivision plays a central role in the analysis of the solvability of tasks in different computational models in computer science (see [2] for details). Simplicial neural networks (SNN) have emerged as a new direction in graph learning, extending the concept of convolutional architectures from node space to simplicial complexes, as detailed in [4, 3, 12]. The reconstruction of random simplices (see [1, 9, 8]) has emerged as a recent and significant trend in the study of random structures, gaining attention for its theoretical importance and potential applications.
@@ 1.2 Our Contributions
The concept of given weights associated to edges in a graph has immense real-life application. With this thought, in this article, we define the concepts of a weighted simplicial complex and a $d$-path in the complex. The previously described real-life example (see Section 1.1) showcases the importance of our constructions in real-life. How to find the shortest path in a weighted simplicial complex is the question that intuitively follows. We provide a novel algorithm to find the shortest path in a weighted simplicial complex. Our algorithm is based on the same underlying principles as Dijkstra’s algorithm for graphs. We also provide correctness arguments for our algorithm along with thorough time and memory complexity analysis.
@@ 1.3 Outline
In Section 2, we provide all the notations and definitions necessary for explaining our algorithm. This section also introduces the concepts of weighted simplicial complex and $d$-path. The section ends with a description of the Dijkstra’s algorithm for graphs. In Section 3 we describe our novel algorithm to find the shortest path in a simplicial complex. The section ends with a detailed analysis of the algorithm. We conclude our study with some remarks and possible future research directions in Section 4.
@@ 2 Preliminaries
In this section, the notations and definitions that will form the building blocks for our algorithm will be discussed. This section also covers all the newly introduced concepts in this paper. Furthermore, for didactic reasons, we also explain Dijkstra’s algorithm for graphs at the end of this section. This will help the reader to relate our set-up with that of the graphs and see how our algorithm can be thought of as a higher-dimensional generalisation of Dijkstra’s algorithm for graphs.
@@ 2.1 The Set-up
Definition 1 (Finite Abstract Simplicial Complex)
@@
For any finite set $A$, let $|A|$ denote the number of elements in $A$. A set $\sigma\in X$ with $|\sigma|=k+1$ is called a $k$-dimensional simplex (or simply a $k$-simplex), where $k$ is a non-negative integer. Specifically, a vertex is a 00-simplex, an edge is a $1$-simplex, a triangle is a $2$-simplex, and so on.
@@
where $\mathcal{X}_{0}:=\{1,2,\dots,n\}$, $X_{k}:=\{\{i_{0},\dots,i_{k}\}\mid 1\leq i_{0}<\dots<i_{k}\leq n\}$ for $1\leq k\leq d$, and $X^{d}\subseteq X_{d}$. Here, $X_{k}$ represents the set of all $k$-simplices on $\mathcal{X}_{0}$. In this model, $X$ is a complex with a complete $((d-1))$-dimensional skeleton.
Definition 2 (Weighted $d$-complex)
@@
We say that $\sigma,\sigma^{\prime}\in X_{(d-1)}$ are neighbors if $\sigma\cup\sigma^{\prime}\in X^{d}$, in which case we write $\sigma\sim\sigma^{\prime}$ (see Figure 2111Given a $d$-dimensional simplicial complex $X$, the figures illustrating $X$ display a selection of $(d-1)$-dimensional faces, rather than the exhaustive set of all $(d-1)$-simplices contained within $X$, in order to maintain visual concision and avoid clutter.). This concept is similar to the one introduced in [10].
@@
Definition 3 ($d$-path and $d$-cycle)
A $d$-path in a $d$-complex $X$ is a sequence of distinct $(d-1)$-simplices $\sigma_{0},\sigma_{1},\dots,\sigma_{l}$ such that $(\sigma_{i}\cup\sigma_{i+1})\in X^{d}$ for all $0\leq i<l$ and the collection $\{\sigma_{i}\cup\sigma_{i+1}:0\leqslant i\leqslant l\}$ is a disjoint collection. The $d$-path is said to have length $l$ and is denoted by $P=(\sigma_{0},\sigma_{1},\dots,\sigma_{l})$. If the $d$-path starts and ends at the same $(d-1)$-simplex, it is called a $d$-cycle of length $l$.
@@
In a weighted $d$-complex $(X,w)$, the shortest $d$-path between two vertices $\sigma_{1}$ and $\sigma_{2}$ is a $d$-path from $\sigma_{1}$ to $\sigma_{2}$ that minimizes the sum of the weights of the $d$-simplices along the $d$-path. The shortest $d$-path may not be unique. The weighted distance between two $(d-1)$-simplices $\sigma_{1},\sigma_{2}\in X_{(d-1)}$ is the sum of the weights of the $d$-simplices along the shortest $d$-path from $\sigma_{1}$ to $\sigma_{2}$. The weighted distance is denoted as $d_{w}(\sigma_{1},\sigma_{2})$.
2.2 Dijkstra’s Algorithm for Graphs
In this section, we briefly recall Dijkstra’s algorithm for graphs. Dijkstra’s algorithm [[11], Algorithm 2.3.5] finds the shortest path from a source $(d-1)$-simplex $\sigma_{0}$ (vertex) to all other vertices in a weighted graph $G=(V,E,w)$. The algorithm iteratively selects the $(d-1)$-simplex with the smallest known distance, relaxes its neighbours, and updates the shortest known distances. On a high level, the algorithm follows the following steps (1-3).
  1. 1.
    Initialize the distance to the source $d(s)=0$ and all other vertices $d(v)=\infty$.
  2. 2.
    Select the unvisited $(d-1)$-simplex $u$ with the smallest $d(u)$, and update the distances to its neighbors:
    $d(v)=\min(d(v),d(u)+w(u,v))$
  3. 3.
    Repeat until all vertices have been visited or the smallest distance is infinity.
3 Our Algorithm for Finding the Shortest Paths in a Weighted Simplicial Complex
@@
In this section, we describe and analyse our novel algorithm. We also provide example of the workings of our algorithm.
@@
Our algorithm is a greedy algorithm to find the shortest $d$-paths from a source $(d-1)$-simplex $\sigma_{0}$ to all other $(d-1)$-simplexes in a weighted complex $(X,w)$, where $w:X_{d}\to\mathbb{R}$ represents the weight function. Note that the weights have to be strictly positive for this algorithm to work. The algorithm works by iteratively selecting the $(d-1)$-simplexes with the smallest known distance and updating the distances to their neighbouring $(d-1)$-simplexes. The pseudo-code of our algorithm is given in Algorithm 1. But for ease of understanding, the high-level idea of our algorithm is described in the following steps.
@@
    Input: A weighted $d$-complex with nonnegative weight function $w$ and a starting $(d-1)$-simplex $\sigma_{0}$. For any $x\in\mathcal{X}_{0}$, the weight of the $d$-simplex $\sigma_{0}\cup\{x\}$ is $w\left(\sigma_{0}\cup\{x\}\right)$; Let $w\left(\sigma_{0}\cup\{x\}\right)=\infty$ if $\sigma_{0}\cup\{x\}\notin X^{d}$.
@@
      Set the distance to the source $(d-1)$-simplex $\sigma_{0}$ as $t(\sigma_{0})=0$.
@@
      Set the distance to all other $(d-1)$-simplexes $\sigma\in X_{(d-1)}\setminus\{\sigma_{0}\}$ as $t(\sigma)=\infty$.
@@
      Select the unvisited $(d-1)$-simplex $\sigma_{1}$ with the smallest known distance $t(\sigma_{1})$. The choice of $\sigma_{1}$ may be more than one.
@@
      Mark $\sigma_{2}$ as visited. If all $(d-1)$-simplexes have been visited or the smallest known distance among the unvisited $(d-1)$-simplexes is infinity, the algorithm terminates.
@@
      The shortest $d$-path from $\sigma_{0}$ to each $(d-1)$-simplex is now known.
Input: A weighted $d$-complex with non-negative weight function $w$ and a starting $(d-1)$-simplex $\sigma_{0}$
Output: The shortest $d$-path from $\sigma_{0}$ to each $(d-1)$-simplex
1 Set the distance to the source $(d-1)$-simplex $\sigma_{0}$ as $t(\sigma_{0})=0$
2 Set the distance to all other $(d-1)$-simplexes $\sigma\in X_{(d-1)}\setminus\{\sigma_{0}\}$ as $t(\sigma)=\infty$

### v2

@@ Abstract
Simplicial complexes are extensively studied in the field of algebraic topology. They have gained attention in recent time due to their applications in fields like theoretical distributed computing and simplicial neural networks. Graphs are mono-dimensional simplicial complex. Graph theory has application in topics like theoretical computer science, operations research, bioinformatics and social sciences. This makes it natural to try to adapt graph-theoretic results for simplicial complexes, which can model more intricate and detailed structures appearing in real-world systems. Though seemingly obvious, we did not find any previous work that looked into this prospect of simplicial complexes.
In this article, we define the concept of weighted simplicial complex and $d$-path in a simplicial complex. Both these concepts have the potential to have numerous real-life applications. We start by adapting the Depth-First Search and Breadth-First Search algorithms for our setup. Next, we provide two novel algorithms to find the shortest paths in a weighted simplicial complex. The core principles of our algorithms align with those of Dijkstra’s algorithm and Bellman-Ford algorithm for graphs. Hence, this work lays a building block for the sake of integrating graph-theoretic concepts with abstract simplicial complexes.
@@
Weighted Simplicial Complex, Shortest Path, Graph Discovery Algorithms, Dijkstra’s Algorithm, Bellman-Ford Algorithm
@@ 1 Introduction
Now a graph can be seen as a boiled down version of a simplicial complex. While a graph is typically a 1-dimensional object, an abstract simplicial complex can exist in any finite dimension. One of the main differences between a graph and a simplicial complex is that in a simplicial complex with $d\geq 2$, more than two (d1)(d-1)-dimensional simplices can be connected by a single $d$-dimensional simplex, whereas in a graph, an edge can connect at most two vertices.
@@
On the other hand, Simplicial complexes have also been used in various fields as detailed in [Herlihy99, Herlihy14, Koz12]. For example, in [Herlihy99] and [Herlihy14], the results on simplicial complexes were used to prove new results in theoretical distributed computing. Distributed computing is a field of computer science that studies distributed systems. In [Koz12], Kozlov studied the chromatic subdivision of a simplicial complex. Chromatic subdivision plays a central role in the analysis of the solvability of tasks in different computational models in computer science (see [attiya2004distributed] for details). Simplicial neural networks (SNN) have emerged as a new direction in graph learning, extending the concept of convolutional architectures from node space to simplicial complexes, as detailed in [ebli2020simplicial, chen2022bscnets, wu2023simplicial]. The reconstruction of random simplices (see [adhikari2022shotgun, ophelders2025sweeping, ma2025reconstructing]) has emerged as a recent and significant trend in the study of random structures, gaining attention for its theoretical importance and potential applications.
@@ 1.2 Our Contributions
The concept of assigning weights to edges in a graph has immense practical importance, forming the basis of many optimization problems in real-world systems. Motivated by this, in this article we introduce two fundamental generalizations to the higher-dimensional setting: (i) the notion of a weighted simplicial complex, and (ii) the definition of a $d$-path in such a complex. Both concepts provide the groundwork for extending classical graph-theoretic algorithms to simplicial complexes, enabling the treatment of higher-dimensional adjacency relations in a combinatorial framework. To the best of our knowledge and much to our astonishment, these concepts have not been defined in the literature before.
Building upon these foundations, our main contributions can be summarized as follows:
  1. 1.
    Generalized Discovery Algorithms: We extend the classical Breadth-First Search (BFS) and Depth-First Search (DFS) algorithms from graphs to simplicial complexes. These algorithms, presented as Algorithm 1 and Algorithm 2, can be thought of as the warm-up for the logical thought chain to our later more intricate algorithms. We also prove their correctness and analyze their time and memory complexities, both of which remain linear in the size of the complex-matching the optimal asymptotic bounds known for graphs.
  2. 2.
    Shortest-Path Algorithm in Weighted $d$-Complexes: We propose a novel greedy algorithm (see Algorithm 3) for computing the shortest $d$-paths in a weighted simplicial complex. Our algorithm generalizes Dijkstra’s algorithm to the simplicial setting while retaining its essential principles of distance relaxation and greedy vertex selection. We rigorously prove its correctness and derive its time and memory complexities, showing that the algorithm runs in $\mathcal{O}((n+m)\log n)$, where $n$ and $m$ denote the numbers of (d1)(d-1)- and $d$-simplices respectively.
  3. 3.
    Relaxation-Based Generalization of Bellman-Ford: Extending beyond non-negative weights, we introduce Algorithm 4, a Bellman-Ford-type procedure that correctly computes shortest weighted (d1)(d-1)-paths even in the presence of negative weights. This algorithm also detects negative $d$-cycles, thus fully generalizing the Bellman-Ford framework to higher-dimensional combinatorial structures.
This work bridges the gap between classical graph algorithms and higher-dimensional combinatorial structures. This offers theoretical depth and potential for real-world applications. Our constructions illustrate how simplicial complexes can serve not only as topological objects but also as computational and algorithmic entities in discrete mathematics and network science. We also give an example where our set-up can model the problem in hand in a better way than usual graphs.
@@ 1.3 Outline
In Section 2, we provide all the notations and definitions necessary for explaining our algorithm. This section also introduces the concepts of weighted simplicial complex and $d$-path. The Section 3 includes the description and analysis of BFS and DFS for our setting. In Section 4 and Section 5 we describe our novel algorithms to find the shortest path in a simplicial complex. We provide a real-life example of our setup in Section 6 that shows how our setup can model some of the real-life problems better than graphs. We conclude our study with some remarks and possible future research directions in Section 7.
@@ 2 Preliminaries
In this section, the notations and definitions that will form the building blocks for our algorithm will be discussed. This section also covers all the newly introduced concepts in this paper.
@@ 2.1 The Set-up
Definition 1 (Finite Abstract Simplicial Complex).
@@
For any finite set $A$, let $|A|$ denote the number of elements in $A$. A set $\sigma\in X$ with $|\sigma|=k+1$ is called a $k$-dimensional simplex (or simply a $k$-simplex), where $k$ is a non-negative integer. Specifically, a vertex is a 0-simplex, an edge is a $1$-simplex, a triangle is a $2$-simplex, and so on.
@@
where $\mathcal{X}_{0}:=\{1,2,\dots,n\}$, $X_{k}:=\{\{i_{0},\dots,i_{k}\}\mid 1\leq i_{0}<\dots<i_{k}\leq n\}$ for $1\leq k\leq d$, and $X^{d}\subseteq X_{d}$. Here, $X_{k}$ represents the set of all $k$-simplices on $\mathcal{X}_{0}$. In this model, $X$ is a complex with a complete ((d1))((d-1))-dimensional skeleton.
Definition 2 (Weighted $d$-complex).
@@
We say that $\sigma,\sigma^{\prime}\in X_{(d-1)}$ are neighbors if $\sigma\cup\sigma^{\prime}\in X^{d}$, in which case we write $\sigma\sim\sigma^{\prime}$ (see Figure 2111Given a $d$-dimensional simplicial complex $X$, the figures illustrating $X$ display a selection of (d1)(d-1)-dimensional faces, rather than the exhaustive set of all (d1)(d-1)-simplices contained within $X$, in order to maintain visual concision and avoid clutter.). This concept is similar to the one introduced in [PRR2017].
@@
Definition 3 ($d$-path and $d$-cycle).
A $d$-path in a $d$-complex $X$ is a sequence of distinct (d1)(d-1)-simplices $\sigma_{0},\sigma_{1},\dots,\sigma_{l}$ such that $(\sigma_{i}\cup\sigma_{i+1})\in X^{d}$ for all $0\leq i<l$ and the collection $\{\sigma_{i}\cup\sigma_{i+1}:0\leqslant i\leqslant l\}$ is a disjoint collection. The $d$-path is said to have length $l$ and is denoted by $P=(\sigma_{0},\sigma_{1},\dots,\sigma_{l})$. If the $d$-path starts and ends at the same (d1)(d-1)-simplex, it is called a $d$-cycle of length $l$.
@@
Figure 4: Example of a 2-Path in a Complex. Here PP= $(\{1,2\}$$\{2,6\}$$\{5,6\}$$\{5,10\}$$\{10,11\}$$\{11,13\})$ is a 2-path from $\{1,2\}$ to $\{11,13\}$.
In a weighted $d$-complex $(X,w)$, the shortest $d$-path between two vertices $\sigma_{1}$ and $\sigma_{2}$ is a $d$-path from $\sigma_{1}$ to $\sigma_{2}$ that minimizes the sum of the weights of the $d$-simplices along the $d$-path. The shortest $d$-path may not be unique. The weighted distance between two (d1)(d-1)-simplices $\sigma_{1},\sigma_{2}\in X_{(d-1)}$ is the sum of the weights of the $d$-simplices along the shortest $d$-path from $\sigma_{1}$ to $\sigma_{2}$. The weighted distance is denoted as $d_{w}(\sigma_{1},\sigma_{2})$.
3 Discovery Algorithms in Simplicial Complexes
Before discussing shortest-path algorithms, we examine two foundational graph-discovery algorithms generalized to simplicial complexes. These procedures, breadth-first search and depth-first search, form the unweighted combinatorial basis of all subsequent weighted algorithms.
Adjacency Graph of (d1)(d-1)-Simplices
Let $(X,w)$ be a weighted $d$-complex with a complete (d1)(d-1)-skeleton. We define the adjacency graph associated with $X$ as
$G(X):=(V,E),\quad V:=X_{d-1},\qquad E:=\{(\sigma,\sigma^{\prime})\in V\times V\mid\sigma\cup\sigma^{\prime}\in X_{d}\}.$
Each node in $G(X)$ corresponds to a (d1)(d-1)-simplex, and an edge connects two nodes if they jointly form a $d$-simplex of $X$. When $d=1$, $G(X)$ coincides with the underlying graph itself.
A traversal of $G(X)$ corresponds to the discovery of (d1)(d-1)-simplices reachable under adjacency; such traversals preserve the local combinatorial structure of the complex and establish its connectivity at the (d1)(d-1)-level.
3.1 Breadth-First Search (BFS) in a $d$-Complex
Breadth-first search explores $G(X)$ level by level, discovering all (d1)(d-1)-simplices at minimal unweighted $d$-path distance from a given starting simplex $\sigma_{0}\in X_{d-1}$.
Input : Connected $d$-complex $X=(X_{d-1},X_{d})$, starting simplex $\sigma_{0}\in X_{d-1}$
Output : Traversal order and distance levels of (d1)(d-1)-simplices from $\sigma_{0}$
1
2Mark all $\sigma\in X_{d-1}$ as unvisited;
3 Initialize an empty queue $Q$;
4 $t(\sigma)\leftarrow+\infty$ for all $\sigma$,  $t(\sigma_{0})\leftarrow 0$;
5 Enqueue $\sigma_{0}$ and mark it visited;
6 while $Q$ not empty do
7 $\sigma\leftarrow$ Dequeue($Q$);
8 foreach $\sigma^{\prime}\in S_{\sigma}$ with $\sigma\cup\sigma^{\prime}\in X_{d}$ do
9    if $\sigma^{\prime}$ unvisited then
10         Mark $\sigma^{\prime}$ as visited;
11       $t(\sigma^{\prime})\leftarrow t(\sigma)+1$;
@@
13         Enqueue $\sigma^{\prime}$;
14      end if
15
16   end foreach
17
18 end while
Algorithm 1 Breadth-First Search (BFS) in a $d$-Complex
Lemma 1 (Correctness of BFS in $d$-Complexes).
Let $X$ be a connected $d$-complex and $\sigma_{0}\in X_{d-1}$. After Algorithm 1 terminates, the value $t(\sigma)$ equals the length of the shortest unweighted $d$-path from $\sigma_{0}$ to $\sigma$, for all $\sigma\in X_{d-1}$. Moreover, the set of parent relations forms a shortest $d$-path tree rooted at $\sigma_{0}$.
Proof.
We proceed by induction on the number of layers discovered. Initially, $t(\sigma_{0})=0$. Suppose the claim holds for all simplices whose distance from $\sigma_{0}$ is at most $k$. When processing such a simplex $\sigma$, each unvisited neighbor $\sigma^{\prime}$ with $\sigma\cup\sigma^{\prime}\in X_{d}$ is assigned distance $t(\sigma^{\prime})=t(\sigma)+1=k+1$. Since every $d$-path to $\sigma^{\prime}$ must pass through some $\sigma$ at distance $k$, this value is minimal. The FIFO queue ensures that simplices are processed in nondecreasing order of $t(\cdot)$, hence the invariant is maintained. Upon termination, $t(\sigma)$ equals the minimal number of $d$-simplices connecting $\sigma_{0}$ and $\sigma$, establishing the result. ∎
Lemma 2 (Complexity of BFS).
Let $N:=|X_{d-1}|$ and $M:=|X_{d}|$. Algorithm 1 runs in time $\mathcal{O}(N+M)$ and requires space $\mathcal{O}(N+M)$.
Proof.
Each (d1)(d-1)-simplex is enqueued and dequeued at most once. Each $d$-simplex is inspected only through its incident (d1)(d-1)-simplices, implying that the total number of adjacency checks is linear in $N+M$. ∎
3.2 Depth-First Search (DFS) in a $d$-Complex
Depth-first search performs a recursive traversal of $G(X)$, producing a spanning forest that captures the hierarchical structure of adjacency among (d1)(d-1)-simplices. This recursive process also serves as the basis for cycle detection and enumeration.
Input : Connected $d$-complex $X=(X_{d-1},X_{d})$
Output : DFS forest and parent relation on $X_{d-1}$
1
2Mark all (d1)(d-1)-simplices as unvisited;
3 foreach unvisited $\sigma\in X_{d-1}$ do
4   DFS($\sigma$);
5 end foreach
6
7Procedure DFS($\sigma$):
8 Mark $\sigma$ as visited;
9 foreach $\sigma^{\prime}\in S_{\sigma}$ with $\sigma\cup\sigma^{\prime}\in X_{d}$ do
10 if $\sigma^{\prime}$ unvisited then
[…]
12   end if
13
14 end foreach
Algorithm 2 Depth-First Search (DFS) in a $d$-Complex
Lemma 3 (DFS Visits and Forest Structure).
Let $X$ be a connected $d$-complex. Algorithm 2 visits each (d1)(d-1)-simplex exactly once and produces a spanning tree of the adjacency graph $G(X)$.
Proof.
Marking ensures that each simplex $\sigma$ is visited exactly once, since recursive calls are made only on unvisited simplices. Parent pointers define a tree structure with unique predecessors except the root. If $X$ has multiple connected components, the outer loop constructs a spanning forest. ∎
Lemma 4 (Complexity of DFS).
Let $N=|X_{d-1}|$ and $M=|X_{d}|$. Algorithm 2 runs in time $\mathcal{O}(N+M)$ and space $\mathcal{O}(N+M)$.
Proof.
Each adjacency $(\sigma,\sigma^{\prime})$ is explored at most once. Recursive depth never exceeds $N$, and the storage for visited markers and parent pointers is linear in $N+M$. ∎
3.3 Essence of Discovery Algorithms
Both BFS and DFS operate solely on the adjacency relation induced by $X_{d}$. They generalize classical graph-traversal schemes to the simplicial setting, discovering the structure of (d1)(d-1)-simplices and their interconnections via $d$-simplices. BFS produces minimal unweighted distances (Lemma 1) and provides a natural initialization for weighted algorithms such as Dijkstra and Bellman-Ford, which extend the scalar distance updates to real weights $w(\sigma\cup\sigma^{\prime})$. DFS, besides its use in connectivity and topological sorting, underlies the cycle-detection and enumeration techniques. Both algorithms have linear complexity in the size of the complex, matching the optimal asymptotic bounds known for graph traversals.
The BFS and DFS algorithms presented above provide the essential discrete infrastructure for subsequent metric computations on simplicial complexes. They yield both structural information (connectivity, spanning trees, and cycles) and the traversal order required by shortest-path procedures, which are examined in the next section.
4 Shortest Paths in Absence of Negative Weights
@@
In this section, we describe and analyze our first algorithm for finding shortest path. We also provide example of the workings of our algorithm.
@@
Our algorithm is a greedy algorithm to find the shortest $d$-paths from a source (d1)(d-1)-simplex $\sigma_{0}$ to all other (d1)(d-1)-simplexes in a weighted complex $(X,w)$, where $w:X_{d}\to\mathbb{R}$ represents the weight function. Note that the weights have to be non-negative for this algorithm to work. The algorithm works by iteratively selecting the (d1)(d-1)-simplexes with the smallest known distance and updating the distances to their neighbouring (d1)(d-1)-simplexes. The pseudo-code of our algorithm is given in Algorithm 3. But for ease of understanding, the high-level idea of our algorithm is described in the following steps.
@@
    Input: A weighted $d$-complex with non-negative weight function $w$ and a starting (d1)(d-1)-simplex $\sigma_{0}$. For any $x\in\mathcal{X}_{0}$, the weight of the $d$-simplex $\sigma_{0}\cup\{x\}$ is $w\left(\sigma_{0}\cup\{x\}\right)$; Let $w\left(\sigma_{0}\cup\{x\}\right)=\infty$ if $\sigma_{0}\cup\{x\}\notin X^{d}$.
@@
      Set the distance to the source (d1)(d-1)-simplex $\sigma_{0}$ as $t(\sigma_{0})=0$.
@@
      Set the distance to all other (d1)(d-1)-simplexes $\sigma\in X_{(d-1)}\setminus\{\sigma_{0}\}$ as $t(\sigma)=\infty$.
@@
      Select the unvisited (d1)(d-1)-simplex $\sigma_{1}$ with the smallest known distance $t(\sigma_{1})$. The choice of $\sigma_{1}$ may be more than one.
@@
      Mark $\sigma_{2}$ as visited. If all (d1)(d-1)-simplexes have been visited or the smallest known distance among the unvisited (d1)(d-1)-simplexes is infinity, the algorithm terminates.
@@
      The shortest $d$-path from $\sigma_{0}$ to each (d1)(d-1)-simplex is now known.
Input: A weighted $d$-complex with non-negative weight function $w$ and a starting (d1)(d-1)-simplex $\sigma_{0}$
Output: The shortest $d$-path from $\sigma_{0}$ to each (d1)(d-1)-simplex
1 Set the distance to the source (d1)(d-1)-simplex $\sigma_{0}$ as $t(\sigma_{0})=0$
2 Set the distance to all other (d1)(d-1)-simplexes $\sigma\in X_{(d-1)}\setminus\{\sigma_{0}\}$ as $t(\sigma)=\infty$